OpenStudy (anonymous):

the last sequence I need to solve a puzzel is this one. there is a logic to it but I can't find it... 992-101-496-55-X-32-124-16

jhonyy9 (jhonyy9):

992/2=496/2=248=x 101+9=110/2=55+9=64/2=32